# Heterogeneous Graph Neural Networks for End-to-End Traffic Assignment and Traffic Flow Learning

### v2

@@ Abstract
The traffic assignment problem is one of the significant components of traffic flow analysis for which various solution approaches have been proposed. However, deploying these approaches for large-scale networks poses significant challenges. In this paper, we leverage the power of heterogeneous graph neural networks to propose a novel data-driven approach for end-to-end traffic assignment and traffic flow learning. Our model integrates an adaptive graph attention mechanism with auxiliary "virtual" links connecting origin-destination node pairs, This integration enables the model to capture spatial traffic patterns across different links, By incorporating the node-based flow conservation law into the overall loss function, the model ensures the prediction results in compliance with flow conservation principles, resulting in highly accurate predictions for both link flow and flow-capacity ratios. We present numerical experiments on urban transportation networks and show that the proposed heterogeneous graph neural network model outperforms other conventional neural network models in terms of convergence rate and prediction accuracy. Notably, by introducing two different training strategies, the proposed heterogeneous graph neural network model can also be generalized to different network topologies. This approach offers a promising solution for complex traffic flow analysis and prediction, enhancing our understanding and management of a wide range of transportation systems.
@@
traffic assignment problem , graph neural network , traffic flow prediction , flow conservation , heterogeneous graph
journal: Transportation Research Part C: Emerging Technologies
@@ 1 Introduction
Traffic assignment plays a significant role in transportation network analysis. Solving the traffic assignment problem (TAP) enables a deeper understanding of traffic flow patterns and provides insights into traffic congestion (Nie et al., 2004). Traffic congestion is a major concern in urban areas, leading to increased travel times, fuel consumption, and environmental pollution. By identifying the bottlenecks in transportation networks, city planners can make a strategic plan to increase road capacities to mitigate traffic congestion. To this end, the primary objective of the traffic assignment problem is to determine the traffic flow distribution and identify traffic bottlenecks on a given road network. For solving traffic assignment problems, there are two major formulations with different assumptions: (1) the user equilibrium (UE) assignment and (2) the system optimum (SO) assignment. The UE assignment identifies an equilibrium state in which drivers between each origin-destination (OD) pair cannot reduce travel costs by unilaterally shifting to another route (Kuang et al., 2021). On the other hand, in the SO assignment, drivers choose the routes collaboratively so that the total system travel time is minimized (Seliverstov et al., 2017).
@@
Traffic assignment problems have appeared in other transportation engineering problems. For instance, city planners leverage traffic assignment to evaluate the network performance under extreme event scenarios, including hurricanes (Zou and Chen, 2020) and earthquake (Silva-Lopez et al., 2022; Liu and Meidani, 2023a). These evaluations ensure the transportation infrastructure aligns with the city’s evolving needs and requirements. Furthermore, the UE-TAP is employed in road network design and retrofit investment optimization as a lower-level constraint in a bi-level optimization problem (Hu et al., 2021; Madadi and de Almeida Correia, 2024).
As a critical preliminary step in setting up the TAPs, it is essential to first obtain accurate origin-destination demands that reflect the realistic travel demands. Multiple research has been proposed to estimate regional OD demands (Zhang et al., 2020; Tang et al., 2021; Zhang et al., 2021), leveraging data from various sensors. Despite these methodological advancements, practical challenges persist. Specifically, sensor failures and malfunctions can impose a significant loss of relevant and reliable OD demand information (Sun et al., 2022). Consequently, the inaccurate estimation of OD demand will hinder the accurate calculation of traffic flows. This issue underscores a research gap in effectively solving TAPs and estimating traffic flows under inaccurate OD demand scenarios. In order to solve TAPs, various efficient algorithms have been proposed (Nie et al., 2004; Di Lorenzo et al., 2015). Despite these advances, it is still computationally demanding to conduct a robust network planning for large-scale networks, when multiple relevant scenarios are considered. This highlights the need to develop scalable solution approaches for TAPs to accommodate diverse network scenarios. In recent years, data-driven approaches have leveraged the availability of large-scale transportation data in various transportation problems. For instance, Ye et al. (2019) combines convolutional neural networks to predict multiple transportation demands using real-world taxi and shared bike demand data. Furthermore, recurrent neural networks (RNNs) have been applied to different prediction tasks in transportation engineering (Zhaowei et al., 2020; Fang et al., 2022). Also, Liu et al. (2023) proposed an implicit neural network to travel choice preferences and equilibrium state modeling from multi-day link flow observations.
@@
Recently, graph neural networks (GNNs) have emerged as a powerful tool in transportation performance analysis since they propagate the node features through a graph structure. Compared with RNNs and CNNs, a key advantage of GNNs is their ability to capture spatial and relational information inherent in transportation networks. Among the recent works, Nishi et al. (2018) developed a traffic signal control algorithm based on graph convolutional networks and reinforcement learning. Liu and Meidani (2022) utilizes a graph convolutional network to analyze the node-to-node accessibility of regional transportation systems under earthquakes. Despite these advancements in neural networks, the literature on applying neural networks to TAPs remains relatively sparse. Fan et al. (2023) integrated CNNs with RNNs for predicting the traffic flow distribution by considering the transportation network as the grid map. Furthermore, Rahman and Hasan (2023) utilizes the graph convolutional network for solving UE-TAP and predicting the traffic flow pattern. However, the aforementioned model encountered a few limitations. First of all, one major limitation of the grid-based CNN model (Fan et al., 2023) is that it cannot fully capture the topologies of the transportation system. Furthermore, these models don’t adequately consider the transportation network under various scenarios, e.g., link capacity reduction due to traffic accidents or lane closures due to maintenance. Besides, they don’t fully investigate the model performance under the situations of incomplete/missing OD demand information. Moreover, how these models perform under out-of-distribution data is not comprehensively explored. These limitations highlight the need for further research to enhance the adaptability and real-world applicability of GNNs for traffic assignment problems.
@@
It should be noted that in the transportation network, a direct link between the origin and destination nodes in general may not exist. This while the traffic flows do in fact depend on the relationship between the OD pairs and OD demands. This observation motivates us to consider introducing the auxiliary links between the origin and destination nodes within the original transportation network, which could encapsulate valuable supplementary information, which is often neglected in the previous literature. In this paper, we propose a novel heterogeneous GNN model for efficient traffic assignment. In addition to roadway links, our model includes virtual links between origin-destination nodes, enriching the model with comprehensive OD demand information and thereby facilitating enhanced feature propagation across the network. Furthermore, we introduce a novel adaptive graph attention mechanism to propagate the node features efficiently. And, the proposed model transforms node embeddings into link embeddings for link flow and flow-capacity ratio estimations. To summarize, the major contributions of this work are as follows: (1) This is the first GNN learning of UE-TAP that integrates interdependencies between origin and destination nodes via a heterogeneous graph structure that consists of physical and virtual links and an adaptive attention-based mechanism; (2) the proposed model is trained using data and the governing conservation law and as such the estimated flows are more accurate; (3) due to the integration of virtual links and regularization based on the conservation law, the performance on unseen graphs is improved. The efficiency and generalization capability of the proposed model are investigated through multiple experiments with different road network topologies, link characteristics, and OD demands.
@@
The remainder of this article is structured as follows. General backgrounds on the traffic assignment problem, the neural network and graph neural network models are presented in Section 2. Section 3 includes the explanation of the proposed heterogeneous graph neural network for traffic assignment and traffic flow learning. Furthermore, the experiments with urban road networks and generalized synthetic networks are presented to demonstrate the accuracy and generalization capability of the proposed framework in Section 4. Finally, the conclusion and discussion of the proposed framework are presented in Section 5.
@@ 2.2 Neural Networks
Without loss of generality, we will start by considering a neural network with only one layer. Given a $p$-dimensional input vector $h^{k}\in\mathbb{R}^{p}$, $q$-dimensional the output $h^{k+1}\in\mathbb{R}^{q}$ of single layer neural network, The single-layer neural network with index $k$ can be expressed as:
@@ 2.3 Graph Neural Network
3 Traffic Assignment and Traffic Flow Learning using Heterogeneous Graph Neural Networks
In this section, we elaborate on the proposed architecture of the heterogeneous graph neural networks for traffic assignment and traffic flow learning. The illustration of the proposed model is shown in Fig. 1. It consists of three modules: graph construction & feature preprocessing module; spatial feature extraction module, and edge prediction module. The detail explanation of each module is described as follows.
Figure 1: The illustration of the heterogeneous graph neural network for traffic flow learning. The proposed model consists of three parts: graph construction & feature preprocessing module; spatial feature extraction module, and edge prediction module. The graph features are first passed into the virtual encoder (V-Encoder) through the virtual links. Then the graph embeddings are passed into the real encoder (R-Encoder) through the real links. The flow-capacity ratio and link flow of each link are calculated using the source node feature, destination node feature, and normalized edge feature.
@@ 3.1 Graph Construction & Feature Preprocessing
The heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{r},\mathcal{E}_{v})$ for traffic flow learning consists of one type of node representing the intersections of road segments and two edge types: real links and virtual links. The real links represent the road segments in the road network, while the virtual links represent the auxiliary link between the origin and destination nodes. The incorporation of the virtual link, as an edge augmentation technique, is strategically employed to facilitate enhanced feature updating. In this model, the node feature attribute is denoted as $\bm{X}_{n}\in\mathbb{R}^{|\mathcal{V}|\times(|\mathcal{V}|+2)}$, where $\mathcal{V}$ represents the graph’s node set. Specifically, each row $\bm{x}_{u}\in\mathbb{R}^{1\times(|\mathcal{V}|+2)}$ within this matrix corresponds to the feature vector of a single node $u\in\mathcal{V}$ including the origin-destination demand as well as geographical coordinates. The feature representation for the entire set of real edges is represented with $\bm{X}_{e,r}\in\mathbb{R}^{|\mathcal{E}_{r}|\times 2}$. Each row in this matrix includes the free-flow travel time and the link capacity for a single link.
@@ 4.1 Experiments on Urban Transportation Networks
The training and testing dataset are obtained by solving UE-TAP with the Frank-Wolfe algorithm (Fukushima, 1984). The algorithm converges when the square root of the sum of the squared differences between the link flow in two successive iterations, normalized by the sum of the values of the link flow, falls below the threshold of 1e-5. The GNN model is implemented using PyTorch (Paszke et al., 2019) and DGL (Wang et al., 2019a). The preprocessing layer consists of a three-layer fully connected neural network with an embedding size of 32. The number of GNN layers in the proposed model is 4, including two V-Encoders and two R-Encoders. The number of heads in the attention block is 8. For hyper-parameter selection, the hidden layer size is chosen as 64, which is common in neural network implementation (Liu and Meidani, 2023c, b). The learning rate and batch size of training are 0.001 and 128, respectively. The weights of $L_{\alpha}$, $L_{f}$, and $L_{c}$ in equation 14 are chosen as 1.0, 0.005, and 0.05, respectively.
We evaluated the performance of our proposed heterogeneous GNN model (referred to by HetGAT) and compared it with three benchmark models: a fully connected neural network (FCNN), a homogeneous graph attention network (GAT), and a homogeneous graph convolution network (GCN). The FCNN consisted of five fully connected layers with an embedding size of 64. The GAT and GCN both have four layers of graph message passing layer, followed by three layers of FCNN with an embedding size of 64. The metrics to evaluate performance include the mean absolute error (MAE), root mean square error (RMSE), and the normalized conversation loss $\tilde{L}_{c}$:
@@
After the training is finished, the model performance is evaluated on the testing set. The experiments under the urban road network are conducted in three different settings. The first setting, referred to as LMH-LMH, involved using all levels of disruption (flow reduction scaling levels) in both the training and testing sets. The second setting, namely L-M, involves training the model on light disruption data and testing it on medium disruption data. The third case, which is labeled as M-H, involves training on medium disruption data and testing on high disruption data. The L-M and M-H scenarios will therefore involve unseen cases that don’t exist in the training data. Figure 5 visualizes the variations in link-wise flow in Sioux Falls, EMA, and Anaheim networks under high disruption. The average coefficients of variation of actual link flow for three networks are 0.297, 0.242, and 0.201, respectively. Figure 6 plotted the predicted value and ground truth of the flow-capacity ratio on multiple samples in the Anaheim network under the LMH-LMH setting. In total, the flow-capacity ratio on 10,000 edges is predicted using HetGAT, GAT, and GCN, respectively. Figure 6 indicates that HetGAT has a relatively higher correlation coefficient, which outperforms GAT and GCN.
@@
Figure 6: Comparison of predicted flow-capacity ratio between ground truth and surrogate modeling in the Anaheim network under LMH-LMH setting.
@@
As an additional experiment, we consider a realistic scenario where the regional OD demand values are incomplete by introducing a random binary mask to the original OD demand, More specifically, given a specific missing ratio, we randomly select a number of OD pairs and mask their corresponding OD demand values as zeros in the input node feature. In this way, the model is expected to learn the inherent patterns and structures of the transportation network, even when some of the demand information is missing. The effectiveness of the proposed model under incomplete OD demand scenarios will be evaluated by comparing the predicted traffic flows against the ground truth data obtained from the complete OD demand. Three missing ratios are considered in the experiment: 20%, 30%, and 40%. The training setting and the hyperparameters remain the same as those in the aforementioned experiments. Table 3 summarizes the results of prediction performance under different missing rate scenarios for the LMH-LMH setting. The FCNN model is not considered in these experiments because of their very poor performance in the previous experiments under full OD demand. HetGAT still outperforms GAT and GCN under different networks and different missing ratios. additionally, the flow predictions by HetGAT have relatively better compliance with the flow conservation law compared with GAT and GCN.

### v3

# End-to-End Heterogeneous Graph Neural Networks for Traffic Assignment

@@ Abstract
The traffic assignment problem is one of the significant components of traffic flow analysis for which various solution approaches have been proposed. However, deploying these approaches for large-scale networks poses significant challenges. In this paper, we leverage the power of heterogeneous graph neural networks to propose a novel end-to-end surrogate model for traffic assignment, specifically user equilibrium traffic assignment problems. Our model integrates an adaptive graph attention mechanism with auxiliary "virtual" links connecting origin-destination node pairs, This integration enables the model to capture spatial traffic patterns across different links, By incorporating the node-based flow conservation law into the overall loss function, the model ensures the prediction results in compliance with flow conservation principles, resulting in highly accurate predictions for both link flow and flow-capacity ratios. We present numerical experiments on urban transportation networks and show that the proposed heterogeneous graph neural network model outperforms other conventional neural network models in terms of convergence rate and prediction accuracy. Notably, by introducing two different training strategies, the proposed heterogeneous graph neural network model can also be generalized to different network topologies. This approach offers a promising solution for complex traffic flow analysis and prediction, enhancing our understanding and management of a wide range of transportation systems.
@@
traffic assignment problem , graph neural network , traffic flow prediction , flow conservation , heterogeneity
@@ 1 Introduction
Traffic assignment plays a significant role in transportation network analysis. Solving the traffic assignment problem (TAP) enables a deeper understanding of traffic flow patterns and provides insights into traffic congestion and environmental impacts (Nie et al., 2004; Cheng and Lin, 2024). The primary objective of the TAP is to determine the traffic flow distribution and identify traffic bottlenecks on a given road network. Such analysis enables city planners to make a strategic plan for mitigating traffic congestion. For solving traffic assignment problems, there are two major formulations with different assumptions (Seliverstov et al., 2017): (1) the user equilibrium (UE) assignment and (2) the system optimum (SO) assignment. The UE assignment identifies an equilibrium state in which drivers between each origin-destination (OD) pair cannot reduce travel costs by unilaterally shifting to another route. On the other hand, in the SO assignment, drivers choose the routes collaboratively so that the total system travel time is minimized.
@@
In addition to network analysis, traffic assignment has broader implications in various research and practical areas such as regional resilience and urban planning. For instance, city planners leverage traffic assignment to evaluate the network reliability under extreme event scenarios, including hurricanes (Zou and Chen, 2020) and earthquake (Liu and Meidani, 2023a). These evaluations ensure the transportation infrastructure aligns with the city’s evolving needs and requirements. Furthermore, the TAPs is employed in road network design and retrofit investment optimization as a lower-level constraint in a bi-level optimization problem (Madadi and de Almeida Correia, 2024).
A preliminary and critical step in setting up TAPs involves obtaining accurate regional OD demands. While there have been significant methodological advancements in OD demand estimation (Tang et al., 2021), the existing practical challenges lead to inaccurate estimation (Sun et al., 2022). To illustrate, the current traffic assignment methodologies don’t fully investigate the model performance under incomplete/missing OD demand information. This issue underscores a research gap in effectively solving TAPs and estimating traffic flows under inaccurate OD demand scenarios.
@@
Neural networks have demonstrated their capability for data imputation and data construction. Recently, convolutional neural networks (CNNs) (Fan et al., 2023) and graph neural networks (GNNs) (Rahman and Hasan, 2023) have been utilized to solve TAPs. Despite their improvement, these models encountered a few limitations. First of all, the CNN-based model cannot fully capture the topologies of transportation systems. Furthermore, the aforementioned models don’t adequately consider the transportation network under various scenarios, e.g., link capacity reduction due to traffic accidents or lane closures due to maintenance. Last but not least, how these models perform under out-of-distribution data is not comprehensively explored. These limitations highlight the need for further research to enhance the adaptability and real-world applicability of neural networks for traffic assignment problems, specifically under the scenarios of unexpected events and incomplete OD demand.
@@
It should be noted that the distribution of link flows depends on the demand levels between different OD pairs. However, generally the graph representation of the transportation system only includes the physical roadway links, and a direct link between origin and destination nodes does not exist in the graph model for OD pairs that are not directly connected. This motivates us to also include virtual links between origin and destination node pairs in addition to the roadway links into the graph model. Considering this, we propose a novel heterogeneous GNN surrogate model, to also integrate comprehensive OD demand information and thereby enhance feature propagation across the network. Furthermore, our proposed model includes a novel adaptive graph attention mechanism to propagate the node features efficiently, a component to transform node embeddings into link embeddings for link flow and flow-capacity ratio estimation, and a conservation-based loss function. To summarize, the major contributions of this work are as follows: (1) This is the first GNN learning of UE-TAP that integrates interdependencies between origin and destination nodes via a heterogeneous graph structure that consists of physical and virtual links and an adaptive attention-based mechanism; (2) the proposed model is trained using data and the governing conservation law and as a result, the estimated flows are more accurate; (3) due to the integration of virtual links and regularization based on the conservation law, the performance on unseen graphs is improved. In this paper, the efficiency and generalization capability of the proposed model are investigated through multiple experiments with different road network topologies, link characteristics, and OD demands.
@@
The remainder of this article is structured as follows. Section 2 provides a review of related literature. General backgrounds on the traffic assignment problem, the neural network and graph neural network models are presented in Section 3. Section 4 includes the explanation of the proposed heterogeneous graph neural network for traffic assignment. Furthermore, the experiments with urban road networks and generalized synthetic networks are presented to demonstrate the accuracy and generalization capability of the proposed framework in Section 5. Finally, the conclusion and discussion of the proposed framework are presented in Section 6.
2 Related Literature
The literature related to the estimation of traffic flows in transportation networks is discussed below under two areas of OD demand estimation and traffic assignment. By drawing insights from these related works, we have developed a robust and comprehensive approach to address the challenges associated with traffic assignment problems, particularly under network disruption events and incomplete OD demand data.
2.1 OD Demand Estimation
The primary challenge in predicting OD demands lies in the inability to directly measure demand using traffic sensors. Instead, researchers have developed models to infer the demand information from aggregated data collected on roadways. For instance, the autoregressive integrated moving average model has been developed to forecast traffic demand across different regions (Deng and Ji, 2011). Also, OD demands were estimated based on high-quality link flow counts using the Kalman filter (Zhou and Mahmassani, 2007). In an optimization framework, Zhang et al. (2020) formulated the OD estimation problem as a quadratic programming and then solved it using the alternating direction method of multipliers.
More recently, neural networks have been widely adopted for OD demand estimation due to their capability to model complex temporal and spatial dependencies in transportation datasets. For instance, Xiong et al. (2020) integrated link graph neural networks with Kalman filters to predict OD demand. Also, Tang et al. (2021) employed a three-dimensional convolution neural network to learn the high-dimensional correlations between local traffic patterns and OD flows. Despite these methodological advancements, practical challenges such as sensor failures and malfunctions pose significant risks. These issues can result in the loss of relevant and reliable OD demand information, leading to inaccuracies in demand estimation, and consequently in the prediction of traffic flows Sun et al. (2022).
2.2 Traffic Assignment Problem
Solving the traffic assignment problem provides a deeper understanding of traffic flow patterns and offers insights into managing traffic congestion. As a result, more efficient and reliable solution of traffic assignment problem has been the focus of many recent studies. For instance, Lee et al. (2003) aimed at improving convergence and computational efficiency in large-scale traffic networks by proposing a conjugate gradient projection to enhance the gradient projection. Furthermore, Babazadeh et al. (2020) proposed a reduced gradient algorithm by selecting non-basic paths, to effectively manage computational complexity. However, these models rely on the assumption that the regional OD demand is accurate. In scenarios where OD demand information is incomplete or missing, conventional approaches tend to underestimate the link-wise traffic flow.
Recently, neural networks have demonstrated remarkable capabilities in data reconstruction and generalization (Zhang et al., 2018; Liu and Meidani, 2024), presenting a potential solution for the issue of missing OD demands. However, the literature on neural networks in TAPs remains relatively scarce. One of the main approaches utilizes convolutional neural networks. For instance, Fan et al. (2023) utilized recurrent CNNs for traffic assignment problems by considering the transportation network as the grid map, specifically under incomplete OD demands. Furthermore, since GNNs are inherently able to capture spatial information from graph topologies (Liu and Meidani, 2022), they were also used to solve TAPs Rahman and Hasan (2023).
2.3 Summary
In summary, despite significant advances enabled by deep learning, a comprehensive graph neural network-based framework for equilibrium-based traffic assignment and link-wise flow estimation remains elusive. Firstly, it is challenging to recover the traffic flow distribution from incomplete OD demand. Secondly, the performance of these models under out-of-distribution data has not been thoroughly investigated. Last but not least, further investigations are required to adequately consider the transportation network under various scenarios, e.g., link capacity reduction and lane closures.
@@ 3.2 Neural Networks
Without loss of generality, we begin by considering a neural network with a single layer. Given a $p$-dimensional input vector $\bm{h}^{k}\in\mathbb{R}^{p}$, $q$-dimensional the output $\bm{h}^{k+1}\in\mathbb{R}^{q}$ of single layer neural network, the layer indexed by $k$ can be expressed as:
@@ 3.3 Graph Neural Network
4 Heterogeneous Graph Neural Networks for Traffic Assignment
In this section, we elaborate on the proposed architecture of the heterogeneous graph neural networks for traffic assignment. The illustration of the proposed model is shown in Fig. 1. It consists of three modules: graph construction & feature preprocessing module; spatial feature extraction module, and edge prediction module. The detail explanation of each module is described as follows.
Figure 1: The illustration of the heterogeneous graph neural network for traffic assignment. The proposed model consists of three parts: graph construction & feature preprocessing module; spatial feature extraction module, and edge prediction module. The graph features are first passed into the virtual encoder (V-Encoder) through the virtual links. Then the graph embeddings are passed into the real encoder (R-Encoder) through the real links. The flow-capacity ratio and link flow of each link are calculated using the source node feature, destination node feature, and normalized edge feature.
@@ 4.1 Graph Construction & Feature Preprocessing
The heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{r},\mathcal{E}_{v})$ for traffic assignment consists of one type of node representing the intersections of road segments and two edge types: real links and virtual links. The real links represent the road segments in the road network, while the virtual links represent the auxiliary link between the origin and destination nodes. The incorporation of the virtual link, as an edge augmentation technique, is strategically employed to facilitate enhanced feature updating. In this model, the node feature attribute is denoted as $\bm{X}_{n}\in\mathbb{R}^{|\mathcal{V}|\times(|\mathcal{V}|+2)}$, where $\mathcal{V}$ represents the graph’s node set. Specifically, each row $\bm{x}_{u}\in\mathbb{R}^{1\times(|\mathcal{V}|+2)}$ within this matrix corresponds to the feature vector of a single node $u\in\mathcal{V}$ including the origin-destination demand as well as geographical coordinates. The feature representation for the entire set of real edges is represented with $\bm{X}_{e,r}\in\mathbb{R}^{|\mathcal{E}_{r}|\times 2}$. Each row in this matrix includes the free-flow travel time and the link capacity for a single link.
@@ 5.1.2 Training setup and model architecture
The training and testing dataset are obtained by solving UE-TAP with the Frank-Wolfe algorithm (Fukushima, 1984). The algorithm converges when the square root of the sum of the squared differences between the link flow in two successive iterations, normalized by the sum of the values of the link flow, falls below the threshold of 1e-5. The GNN model is implemented using PyTorch (Paszke et al., 2019) and DGL (Wang et al., 2019a). The preprocessing layer consists of a three-layer fully connected neural network with an embedding size of 32. The number of GNN layers in the proposed model is 4, including two V-Encoders and two R-Encoders. The number of heads in the attention block is 8. For hyper-parameter selection, the hidden layer size is chosen as 64, which is common in neural network implementation (Liu and Meidani, 2023b). The learning rate and batch size of training are 0.001 and 128, respectively. The weights of $L_{\alpha}$, $L_{f}$, and $L_{c}$ in equation 14 are chosen as 1.0, 0.005, and 0.05, respectively.
We evaluated the performance of our proposed heterogeneous GNN model (referred to by HetGAT) and compared it with three benchmark models: a fully connected neural network (FCNN), a homogeneous graph attention network (GAT), and a homogeneous graph convolution network (GCN) (Rahman and Hasan, 2023). The FCNN consisted of five fully connected layers with an embedding size of 64. The GAT and GCN both have four layers of graph message passing layer, followed by three layers of FCNN with an embedding size of 64. The metrics to evaluate performance include the mean absolute error (MAE), root mean square error (RMSE), and the normalized conservation loss $\tilde{L}_{c}$:
@@ 5.1.3 Numerical results
After the training is finished, the model performance is evaluated on the testing set. The experiments under the urban road network are conducted in three different settings. The first setting, referred to as LMH-LMH, involved using all levels of disruption (flow reduction scaling levels) in both the training and testing sets. The second setting, namely L-M, involves training the model on light disruption data and testing it on medium disruption data. The third case, which is labeled as M-H, involves training on medium disruption data and testing on high disruption data. The L-M and M-H scenarios will therefore involve unseen cases that don’t exist in the training data. Figure 5 shows the variations in link flows over the three networks under high disruption. The average coefficients of variation of actual link flow for three networks are 0.297, 0.242, and 0.201, respectively. Furthermore, Figure 6 shows the predicted value and ground truth of the link flows on multiple samples in the Anaheim network under the LMH-LMH setting. In total, the flow-capacity ratio on 10,000 edges is predicted using HetGAT, GAT, and GCN, respectively. Also, to compare the predictions with ground truth, ccorrelation coefficients and pairwise comparisons are shown in Figure 6, indicating that HetGAT outperforms GAT and GCN.
@@
Figure 6: Comparison of link flow between ground truth and surrogate model prediction in the Anaheim network under LMH-LMH setting.
@@
5.1.4 Incomplete OD demand data
As an additional experiment, we consider a realistic scenario where the regional OD demand values are incomplete by introducing a random binary mask to the original OD demand, More specifically, given a specific missing ratio, we randomly select a number of OD pairs and mask their corresponding OD demand values as zeros in the input node feature. In this way, the model is expected to learn the inherent patterns and structures of the transportation network, even when some of the demand information is missing. The effectiveness of the proposed model under incomplete OD demand scenarios will be evaluated by comparing the predicted traffic flows against the ground truth data obtained from the complete OD demand. Three missing ratios are considered in the experiment: 20%, 30%, and 40%. The training setting and the hyperparameters remain the same as those in the aforementioned experiments. Table 3 summarizes the results of prediction performance under different missing rate scenarios for the LMH-LMH setting. The FCNN model is not considered in these experiments because of their very poor performance in the previous experiments under full OD demand. HetGAT still outperforms GAT and GCN under different networks and different missing ratios. additionally, the flow predictions by HetGAT have relatively better compliance with the flow conservation law compared with GAT and GCN. The superior performance of HetGAT in scenarios with missing OD demand can be attributed to inclusion of virtual OD links and an adaptive attention mechanism, which captures the inherent correlation between OD demand features and link flows from a lower-dimensional embedding space. With the help of the V-encoder and R-encoder, HetGAT seeks to align incomplete demand embeddings with complete OD demand embeddings in the embedding space. This alignment enables the model to effectively reconstruct full OD demands from partial data, thereby enhancing the accuracy of its predictions.